### v1

# Enhancing Multimodal Recommendations with Vision-Language Models and Information-Aware Fusion

Abstract.
Recent advances in multimodal recommendation (MMR) have shown that incorporating rich content sources such as images and text can lead to significant gains representation quality. However, existing methods often rely on coarse visual features and uncontrolled fusion, leading to redundant or misaligned representations. As a result, visual encoders often fail to capture salient, item-relevant semantics, limiting their contribution in multimodal fusion. From an information-theoretic perspective, effective fusion should balance the unique, shared, and redundant information across modalities, preserving complementary cues while avoiding correlation bias. This paper presents VLIF, a vision-language and information-theoretic fusion framework that enhances multimodal recommendation through two key components. (i) A VLM-based visual enrichment module generates fine-grained, title-guided descriptions to transform product images into semantically aligned representations. (ii) An information-aware fusion module, inspired by Partial Information Decomposition (PID), disentangles redundant and synergistic signals across modalities for controlled integration. Experiments on three Amazon datasets demonstrate that VLIF consistently outperforms recent multimodal baselines and substantially strengthens the contribution of visual features [^1] Code is available at: https://github.com/dangkh/VLIF.
1. Introduction
Recently, multimodal recommendation (MMR) has emerged as a promising direction by incorporating diverse item-related content, including images, textual descriptions, and structured metadata. These additional sources offer rich semantic cues that complement behavioral signals, for instance, product images often emphasize features designed to attract buyers, providing valuable information.
Figure 1. [Top] Our VLIF outperforms recent SOTA multimodal models. [Bottom] Using VLM to generate visual description of item.
The second focuses on how to fuse heterogeneous modalities to construct expressive and semantically aligned representations (Xu et al., 2025a). Despite these advances, the visual modality has a limited contribution to overall performance (Zhou and Shen, 2023; Xu et al., 2025b), as removing image features typically results in only minor performance drops while textual signals remain dominant (Figure 1[Top]). In practice, product images are often optimized for marketing aesthetics rather than descriptive accuracy, which can mislead visual encoders and reduce the semantic fidelity of learned representations. This limitation highlights the need for incorporating textual context to guide visual understanding. Vision–Language Models (VLMs) offer a promising direction for enhancing multimodal recommendation by providing detailed, semantically aligned representations that capture fine-grained item characteristics and mitigate the limitations of raw visual features (Liu et al., 2024; Zhou et al., 2025). However, generic captioning or summarization often produces off-target content that fails to reflect the true semantics of the item. As illustrated in Figure 1[Bottom], when a product image contains multiple objects (e.g., a mother holding her daughter), a VLM prompted without item context may focus on irrelevant details, such as clothing rather than the actual baby carrier. Incorporating the item’s textual information (e.g., its title) allows the model to generate accurate, context-aware descriptions, demonstrating the value of text-guided visual enrichment in multimodal recommendation. Nevertheless, naive fusion strategies often treat modalities equally (Zhou et al., 2023; Xu et al., 2025b), introducing redundancy, conflicting cues, or over-reliance on text, which weakens representation quality. To overcome this, Partial Information Decomposition (PID) (Liang et al., 2023) provides a principled means to quantify modality interactions, allowing multimodal models to isolate complementary and consistent signals that enhance recommendation performance.
To address these issues, we propose VLIF, a novel framework for multimodal recommendation. Instead of directly relying on raw image features, VLIF leverages vision–language models (VLMs) guided by item information (e.g., titles) to generate fine-grained, context-aware descriptions that capture essential item semantics. Furthermore, an information-aware fusion module inspired by PID is designed to quantify synergistic and redundant interactions between modalities, thereby enhancing multimodal representation learning. To summarize, the main contributions of this work are: First, we propose the first framework that leverage VLM to turn item images into enriched, meaningful features, reducing the limits of noisy or off-target representation. Second, we introduce a fusion module based on PID that strengthen the multimodal fusion. Third, extensive experiments on three Amazon product datasets demonstrate that VLIF consistently outperforms strong multimodal baselines.
Figure 2. Our VLIF pipeline
2. Methodology
We formulate our multimodal recommendation framework VLIF as follows. Let $\mathcal{U}$ and $\mathcal{I}$ denote the user and item sets. Each item $i\in\mathcal{I}$ is associated with a visual feature $\mathbf{v}_{i}$ and a textual feature $\mathbf{t}_{i}$, representing the visual and textual modalities, respectively.
As illustrated in Figure 2, our framework adopts a graph-based architecture that captures both heterogeneous and homogeneous relations to model multimodal user–item interactions. Each item embedding combines textual information with visual features, enhanced through a VLM-based Visual Enrichment module, which generates semantically aligned representations from item images. The resulting multimodal embeddings are further refined by an Information-Aware Fusion module for recommendation.
2.1. Multimodal Recommendation Framework
Previous studies (Zhang et al., 2021; Zhou et al., 2023) show that combining heterogeneous and homogeneous relations improves multimodal recommendation. Based on this insight, we build our framework upon a User–Item graph to model high-order collaborative relations across modalities.
User–Item Graph. We take user embeddings and modality-specific item features as input and construct two user–item graphs $\mathcal{G}=\{\mathcal{G}_{m}\mid g_{v},g_{t}\}$ to model modality-dependent interactions. Each graph $\mathcal{G}_{m}$ maintains an identical topology while preserving only the node features of modality $m$.
Formally, the user and item representations at the $l$-th graph convolution layer are updated as:
(1) $\mathbf{u}_{m}^{(l)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{|\mathcal{N}_{u}|}\sqrt{|\mathcal{N}_{i}|}}\mathbf{i}_{m}^{(l-1)},\quad\mathbf{i}_{m}^{(l)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{|\mathcal{N}_{u}|}\sqrt{|\mathcal{N}_{i}|}}\mathbf{u}_{m}^{(l-1)}.$
Here, $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the neighbor sets of user $u$ and item $i$ in $\mathcal{G}$, respectively. The initial user embedding $\mathbf{u}_{m}^{(0)}$ is randomly initialized, while the initial item embedding $\mathbf{i}_{m}^{(0)}$ is derived from its corresponding visual feature $\mathbf{v}_{i}$ or textual feature $\mathbf{t}_{i}$. The final embedding for each modality is obtained by aggregating features across all $L$ layers:
(2) $\bar{\mathbf{u}}_{m}=\sum_{l=0}^{L}\mathbf{u}_{m}^{(l)},\quad\bar{\mathbf{i}}_{m}=\sum_{l=0}^{L}\mathbf{i}_{m}^{(l)}.$
Each modality-specific user and item embedding is then fed into the Information-Aware Fusion module (Section 2.3) to obtain unified representations $\mathbf{u}_{f}$ and $\mathbf{i}_{f}$, which are subsequently propagated through the homogeneous graphs to capture user–user and item–item relations.
Item–Item Graph. To model semantic correlations between items, we construct an item–item homogeneous graph via a $k$-nearest neighbor (KNN) strategy using item features from each modality $m$. The similarity between items $(i,i^{\prime})\in\mathcal{I}$ is measured by cosine similarity, and the final similarity score is defined as $S_{i,i^{\prime}}=\tfrac{1}{2}(S_{i,i^{\prime}}^{v}+S_{i,i^{\prime}}^{t})$, combining visual and textual similarities. Only the top-$k$ neighbors are retained, while the remaining edges are pruned. The item representation is updated as:
(3) $\mathbf{h}_{i}^{(l)}=\sum_{i^{\prime}\in\mathcal{N}_{i}}S_{i,i^{\prime}}\mathbf{h}_{i^{\prime}}^{(l-1)},$
where $\mathcal{N}_{i}$ denotes the neighbor set of item $i$ in the item–item graph and $\mathbf{h}_{i}^{(0)}$ is initialized with $\mathbf{i}_{f}$. For the User–User Graph, edges are constructed based on the number of co-interacted items, and only the top-$k$ neighbors are preserved. User representations are updated analogously to the aggregation process in Equation (3). The final user and item representations, $\mathbf{z}_{u}$ and $\mathbf{z}_{i}$, are obtained by integrating the homogeneous graph embeddings with the residual fusion signals. The predicted interaction score is then computed as:
(4) $\hat{y}_{ui}=\mathbf{z}_{i}^{\top}\mathbf{z}_{u}.$
2.2. VLM-based Visual Enrichment
We leverage vision–language models (VLMs) to transform visual inputs into semantically enriched textual representations through task-specific prompting. Given a product image $\mathbf{v}_{i}$ and its title $\mathbf{t}^{title}_{i}$, the prompt $\mathcal{P}(\mathbf{v}_{i},\mathbf{t}^{title}_{i})$ integrates both modalities to guide the VLM toward generating item-relevant descriptions.
Figure 3. VLM enriches visual features
To improve semantic reasoning and detail accuracy, we employ a Chain-of-Thought (CoT) prompting strategy that encourages step-by-step generation of enriched item descriptions. As illustrated in Figure 3, the VLM produces descriptive text, which is then encoded into a dense embedding serving as the enhanced visual representation:
(5) $\mathbf{v}_{i}=\mathrm{Enc}\big(\mathrm{VLM}(\mathcal{P}(\mathbf{v}_{i},\mathbf{t}^{title}_{i}))\big),$
where $\mathrm{Enc}(\cdot)$ denotes a text encoder that maps generated descriptions into the feature space.
2.3. Information-Aware Fusion
From the information-theoretic perspective, the relationship of multimodal inputs $X_{i}=\{v_{i},t_{i}\}$ and target $Y$ can be decomposed as:
(6) $I(X_{1},X_{2};Y)=R+S+U_{v}+U_{t},$
where $R$ denotes the redundant information shared by both modalities, $S$ the synergistic information that emerges only through their joint interaction, and $U_{v},U_{t}$ the unique contributions of the visual and textual modalities, respectively. To realize this decomposition in representation space, we explicitly estimate $S$ and $R$, and derive the unique components based on these estimates.
Cross-Modal Synergy Estimation (CSE). We first apply a cross-modal Transformer to capture interactive information between the two modalities:
(7) $\mathbf{h}_{v\rightarrow t}=\mathrm{Trans}(\mathbf{v},\mathbf{t}),$
where $\mathbf{h}_{v\rightarrow t}$ represents attended features obtained by querying vision over text (and symmetrically $\mathbf{h}_{t\rightarrow v}$ for text over vision). The shared information between these two attention directions constitutes the synergistic signal, which we estimate as:
(8) $\mathbf{s}=\tfrac{1}{2}(\mathbf{h}_{v\rightarrow t}+\mathbf{h}_{t\rightarrow v}).$
Transformer-based Redundancy Estimation (TBR). To extract the modality-shared (redundant) information $R$, we construct three input variants:
(9) $\mathbf{X}=[\mathbf{v},\mathbf{t}],\quad\mathbf{X}_{v}=[\mathbf{v},\texttt{[MASK]}_{t}],\quad\mathbf{X}_{t}=[\texttt{[MASK]}_{v},\mathbf{t}],$
where $\texttt{[MASK]}_{t}$ and $\texttt{[MASK]}_{v}$ are special tokens used to remove a modality. Each variant is passed through a shared Transformer encoder $F(\cdot)$ followed by a projection layer $g(\cdot)$:
(10) $\mathbf{h}=g(F(\mathbf{X})),\quad\mathbf{h}_{v}=g(F(\mathbf{X}_{v})),\quad\mathbf{h}_{t}=g(F(\mathbf{X}_{t})).$
Aligning the unimodal encodings $\mathbf{h}_{v}$ and $\mathbf{h}_{t}$ with the full multimodal encoding $\mathbf{h}$ allows the model to capture modality-invariant information, which we define as the redundant feature:
(11) $\mathbf{r}=\tfrac{1}{3}(\mathbf{h}+\mathbf{h}_{v}+\mathbf{h}_{t}).$
Fusion. To obtain the unique visual contribution $U_{v}$, we remove the redundant component $\mathbf{r}$ from the visual representation $\mathbf{v}$ by orthogonal projection:
(12) $\mathbf{v}^{\prime}=\mathbf{v}-\text{Proj}_{\mathbf{r}}(\mathbf{v}),\quad\text{Proj}_{\mathbf{r}}(\mathbf{v})=\frac{\mathbf{v}^{\top}\mathbf{r}}{\|\mathbf{r}\|_{2}^{2}}\mathbf{r}.$
This yields $\mathbf{v}^{\prime}$, the unique and redundancy-free visual representation orthogonal to $\mathbf{r}$. The final item representation concatenates the three complementary components—textual, synergistic, and unique visual features—as:
[…]
where $||$ denotes concatenation. To model user-specific modality preferences, we introduce a learnable weight vector $\textbf{u}^{w}=[u^{w}_{0},u^{w}_{1},u^{w}_{2}]$ derived from user IDs, producing a fused user embedding:
[…]
2.4. Optimization
To regulate modality interactions, we employ InfoNCE that explicitly models the synergy information shared between modalities.
(15) $\mathcal{L}_{\text{s}}=\mathcal{L}_{\text{InfoNCE}(h_{v\rightarrow t},h_{t\rightarrow v})}$
where $\mathcal{L}_{\text{InfoNCE}}$ is the standard contrastive loss that promotes alignment between 2 presentations. Similarly, we align the modality’s common information, through the three representations:
(16) $\mathcal{L}_{\text{r}}=\mathcal{L}_{\text{InfoNCE}}(h,h_{v})+\mathcal{L}_{\text{InfoNCE}}(h,h_{t})+\mathcal{L}_{\text{InfoNCE}}(h_{v},h_{t}),$
which maximizes the mutual information among $(h,h_{v},h_{t})$ and drives the model to capture a modality-invariant redundant representation $\mathbf{r}$. The final training objective is formulated as:
(17) $\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda(\mathcal{L}_{\text{s}}+\mathcal{L}_{\text{r}}).$
where $\mathcal{L}_{\text{rec}}$ is the recommendation BPR loss and $\lambda$ is a hyperparameter controlling the strength of fusion regularization.
3. Experiments
3.1. Experimental Settings
Dataset Metric BPR LightGCN VBPR MMGCN GRCN SLMRec LATTICE FREEDOM DRAGON MENTOR COHESION VLIF Improv.
[…]
Baby R@10 0.0357 0.0479 0.0423 0.0421 0.0532 0.0521 0.0547 0.0627 0.0513 0.0678 0.0680 0.0705 3.7%
R@20 0.0575 0.0754 0.0663 0.0660 0.0824 0.0772 0.0850 0.0992 0.0803 0.1048 0.1052 0.1083 2.9%
N@10 0.0192 0.0257 0.0223 0.0220 0.0282 0.0289 0.0292 0.0330 0.0278 0.0362 0.0354 0.0378 4.4%
N@20 0.0249 0.0328 0.0284 0.0282 0.0358 0.0354 0.0370 0.0424 0.0352 0.0450 0.0454 0.0474 4.4%
Sports R@10 0.0432 0.0569 0.0558 0.0401 0.0599 0.0663 0.0620 0.0717 0.0588 0.0763 0.0752 0.0788 3.3%
R@20 0.0653 0.0864 0.0856 0.0636 0.0919 0.0990 0.0953 0.1089 0.0899 0.1139 0.1137 0.1168 2.5%
N@10 0.0241 0.0311 0.0307 0.0209 0.0330 0.0365 0.0335 0.0385 0.0324 0.0409 0.0409 0.0430 5.1%
N@20 0.0298 0.0387 0.0384 0.0270 0.0413 0.0450 0.0421 0.0481 0.0404 0.0511 0.0503 0.0528 3.3%
Clothing R@10 0.0206 0.0361 0.0281 0.0227 0.0421 0.0442 0.0492 0.0629 0.0452 0.0668 0.0665 0.0681 1.9 %
R@20 0.0303 0.0544 0.0415 0.0361 0.0657 0.0659 0.0733 0.0941 0.0675 0.0989 0.0983 0.1013 2.4%
N@10 0.0114 0.0197 0.0158 0.0120 0.0224 0.0241 0.0268 0.0341 0.0242 0.0360 0.0358 0.0374 3.9%
N@20 0.0138 0.0243 0.0192 0.0154 0.0284 0.0296 0.0330 0.0420 0.0298 0.0441 0.0438 0.0462 4.8%
Table 1. Overall performance. Bold indicates best performance in each row.
Module Approach Baby Sports
R@10 R@20 N@20 R@10 R@20 N@20
Image Feature Original 0.0668 0.1026 0.0450 0.0751 0.1126 0.0510
VLM w.o title 0.0692 0.1077 0.0469 0.0770 0.01159 0.0515
Multimodal Fusion Pooling 0.0668 0.1031 0.0452 0.0758 0.1130 0.0506
Concat 0.0677 0.1035 0.0454 0.0778 0.1142 0.0511
Weighted Concat 0.0668 0.1068 0.0461 0.0770 0.1160 0.0520
VLIF 0.0705 0.1083 0.474 0.0788 0.1168 0.0528
Table 2. Ablation study. Contribution of each sub-module.
Dataset. Following prior studies (Xu et al., 2025b; Zhou and Shen, 2023), we conduct experiments on three categories of the Amazon Review dataset [^2] Publicly available at http://jmcauley.ucsd.edu/data/amazon/links.html : Baby, Sports and Outdoors, and Clothing, Shoes and Jewelry, denoted as Baby, Sports, and Clothing for brevity. We adopt the standard 5-core setting, ensuring that each user and item has at least five interactions.
Evaluation. We compare VLIF with the following SOTA methods: BPR, LightGCN (He et al., 2020), VBPR (He and McAuley, 2016), MMGCN (Wei et al., 2019), GRCN (Wei et al., 2020), SLMRec (Tao et al., 2023), LATTICE (Zhang et al., 2021), FREEDOM (Zhou and Shen, 2023), DRAGON (Zhou et al., 2023),MENTOR (Xu et al., 2025b), COHESION (Xu et al., 2025a). Performances are evaluated on 2 metrics: Recall@$10$ (R@1010) and NDCG@$10$ (N@1010). We follow the setting of  (Zhou and Shen, 2023) with a random data splitting of 8:1:1 for training, validation, and testing.
Implementation Details. We implement VLIF with Xavier initialization and a batch size of 1024. Hyperparameters are set with a learning rate of 0.001 and embedding dimension $d=64$. The model uses two layers for user–item, item–item, and user–user graphs, with top-$k=10$ and $\lambda=0.1$. Image descriptions are generated using the Qwen-VL 7B model. Experiments are conducted on an NVIDIA A5000 (24GB) GPU. VLM-generated descriptions and textual content (concatenated from each item’s brand, title, description, and category) are both encoded using Sentence-BERT [^3] https://huggingface.co/sentence-transformers.
3.2. Experimental Results
Overall Performance. As shown in Table 1, our VLIF consistently outperforms the two strongest baselines, MENTOR and COHESION, by a notable margin across all datasets and evaluation metrics. This demonstrates that integrating features derived from VLMs with our information-aware fusion module yields superior performance compared to using raw features or conventional fusion strategies, providing an effective solution for multimodal representation learning. When compared with (Zhou et al., 2023), which adopts a similar multimodal encoder, replacing the original visual features with VLM-based representations and our fusion strategy further improves performance.
Ablation. Furthermore, we investigate the contribution of each proposed module. Table 2 reports the performance when removing either the VLM-based feature generation or the information-aware fusion. The performance degradation observed highlights the necessity of both components. In particular, for image features, the original visual embeddings yield the lowest accuracy, followed by plain VLM features, while our task-specific VLM guided by item titles (VLIF) achieves the best performance. Likewise, alternative fusion strategies perform worse than our proposed information-aware fusion. Moreover, Figure 1 illustrates that incorporating VLM features balances the contribution of each modality, preventing over-reliance on textual signals. Finally, in Figure 4, we also evaluate VLIF with different VLM backbones, and the results confirm that our framework maintains stable and consistent improvements across models, demonstrating its general applicability. Furthermore, we visualize the original and VLM-based visual embeddings. The original features are scattered with many outliers, while the VLM-based embeddings show a more structured distribution, indicating richer and more semantically aligned representations.
4. Conclusion
In this work, we propose VLIF, a multimodal recommendation framework that leverages VLM to generate fine-grained item representations and an information-aware fusion module to integrate modalities effectively. Experiments on three Amazon product datasets demonstrate that VLIF consistently outperforms strong multimodal baselines and significantly strengthens the contribution of the visual modality. In future work, we plan to further enhance the framework by jointly training the VLM with recommendation signals, enabling task-adaptive visual–textual understanding for recommendation.
Figure 4. Experiments on Baby. [Top] Item embedding visualization. [Bottom] Performance of VLM models.